# Integrability in Perturbed Black Holes: Background Hidden Structures

### v1

@@ II.3 Darboux covariance of BH perturbations
where $\mathcal{C}_{R}$ is an arbitrary integration constant. Notice that logarithmic derivatives of a time independent Schrödinger equation automatically satisfy the Riccati equation, e.g. $W_{0}$ satisfies it with $\mathcal{C}_{R}=k_{0}^{2}$. The consistency between the expressions for $W$ and $W_{,x}$, as found from Eqs. (39) and (40), gives the following equation for the difference $\delta V=V-v\,$
@@ III Integrable structures in Cauchy slices
We have described the complete landscape of master equations describing BH perturbations, i.e. pairs of master functions and potentials. We have also seen the role of the Darboux transformations as a way of connecting all of them and establishing the isospectrality of the underlying physical description corresponding to this infinite set. In this sense, the Darboux transformation plays the role of a covariance or symmetry of this set of possible physical descriptions of (vacuum) BH perturbations, establishing their physical equivalence. These results provide the correct way to separate the slow and fast DoFs in this context. In this section, we show the hidden integrable structure associated to the nonlinear slow DoFs, i.e. the potentials.
@@ III.1.1 Darboux transformation and integrability of the KdV equation
where $\sigma$ is the time of the nonlinear KdV dynamics [^9] Notice that in previous works [14, 15, 16] we used $\tau$ for the KdV evolution instead of $\sigma$. Here we use $\tau$ for the time along the hyperboloidal slicing., which does not have to be confused with the Schwarzschild time $t$. From the point of view of BH physics, $\sigma$ does not have in principle any role. One interesting analogy with the original KdV equation, which was introduced to study the propagation of waves in shallow water, is that in this context $V$ is essentially the approximation to the gravitational Newtonian potential around the Earth surface, while in our case it is going to be exactly any of the BH potentials (see Sec. II.2).
@@ III.1.2 Isospectrality under KdV flows
The KdV equation, as well as other integrable systems, admits a particular bi-Hamiltonian structure, meaning that one can find two different (but compatible) Poisson structures and Hamiltonians such that the KdV equation can be described in terms of two different by analogous Hamiltonian flows [101]. Starting from the existence of infinite conservation laws [29], Gardner was the first to realize that the KdV equation can be cast as an Hamiltonian system [102]. A different approach was taken by Zakharov and Faddeev [30], who combined inverse scattering techniques together with the Hamiltonian formalism to find the action-angle variables of the system in terms of the Bogoliubov coefficients, which completely defines the scattering matrix in the linear Schrödinger system.
@@ III.2 The structural triangle KdV-Virasoro-Schwarzian derivative
We are rewriting them together to highlight the fact that the fundamental structure responsible for both transformations (the conformal transformations and the KdV deformations) are the same and are given by the Magri bracket. Indeed, this is no coincidence as we can consider the KdV deformations as a restricted family of infinitesimal conformal transformations, as in Eq. (87), by taking
@@ III.3 Conformal transformations in the frequency domain
We are going to restrict ourselves to transformations that cancel the linear term in $\tilde{\psi}_{,\tilde{x}}$, so that the final equation resembles the original Schrödinger equation [Eq. (31)]. By looking into Eq. (102), we see that cancelling the term in $\tilde{\psi}_{,\tilde{x}}$ means to impose the following different relation between $\omega$ and $\tilde{x}(x)$:
@@
Therefore, the potential transforms according to the typical conformal transformation of (the holomorphic components of) the energy-momentum tensor of a $2d$ CFT, i.e.
@@
Notice that the transformation of the wave function $\psi$ reminds the conformal transformation, under the mapping $x\mapsto\tilde{x}=\tilde{x}(x)$, of a primary field [110]. Indeed, the KdV/Virasoro structure described in Sec. III.2 relies on the extension to the complex plane (complexification of the independent variable $x$) and on treating the new complex variables $z$ and $\bar{z}$ as independent. The conformal transformation of a primary field in this space is then typically written as
@@ IV Integrable structures in hyperboloidal slices
We have investigated the emergence of an underlying integrable structure associated to the slow DoFs, i.e. the solitonic potential response of the BH to gravitational perturbations, appearing when studying the linear dynamics of the fast DoFs, the perturbations, in Cauchy slices. As we already mentioned in the Introduction, the separation into slow/fast DoFs shows a natural association to bulk/boundary properties of the system once an hyperboloidal foliation of the Lorentzian manifold $M_{2}$ is considered [50]. In addition, the hyperboloidal slicing is particularly meaningful when studying properties of the system at null infinities. In our framework, it represents a powerful tool to connect bulk and boundary properties of the system. Indeed, we introduce in this section the necessary following step in the construction: the preservation of the integrable structure related to the slow DoFs in the hyperboloidal slicing. This sets the stage for future investigations analyzing the interplay with the boundary/asymptotic features of the system.
In this section we first review the typical hyperboloidal slicing for master equations and then show that it is invariant under scale transformations of the master function. This feature, of interest by itself, turns out to be crucial in the study of how the bulk symmetries previously described in the Cauchy slices, leak in the hyperboloidal picture, while consistently preserving the bulk/boundary interpretation.
@@ IV.1 Basics of the hyperboloidal slicing
There is an interesting feature of the structure of this operator which provides a different perspective on the separation in slow and fast DoFs. It was shown in Refs. [50, 41] that the non-selfadjointedness of the operator $\mathbb{L}$ is associated to the operator $\mathcal{L}_{2}$, which account for the dissipative character of the system at ${I}^{+}$ and the horizon ($r=r_{s}$), i.e. the asymptotic boundaries. Taking this into account, the operator $\mathcal{L}_{1}$ is associated to “bulk” properties of the system while $\mathcal{L}_{2}$ to “boundary” properties. This supports the conclusion that, within perturbation theory, the properties of the system are expressed by the slow DoFs, in the sense explained above (see Refs. [50, 41] for more detailed arguments).
@@ IV.2 Scale covariance in hyperboloidal slicing
$\mathbb{\tilde{L}}=\frac{1}{i}\left(\begin{array}[]{cc}0&1\\[5.69054pt] \tilde{\mathcal{L}}_{1}&\tilde{\mathcal{L}}_{2}\end{array}\right)\,,$ (162)
@@
This shows the complete covariance of the hyperboloidal approach under transformations of the master function by a scale factor, as indeed the $\mathcal{L}$ operator preserves the same structure as in the ordinary case (135), with the operator ~$\tilde{\mathcal{L}}_{1}$ having a Sturm-Liouville form. This non-trivial result introduces an apparently arbitrary freedom which may turn out to be useful in several application, including numerical calculations.
@@ IV.3 Conformal hyperboloidal slicing
$-\left[1-\left(\frac{h^{\prime}}{g^{\prime}}\right)^{2}\right]\partial^{2}_{\sigma}\tilde{\phi}+\frac{2h^{\prime}}{(g^{\prime})^{2}}\partial^{2}_{\sigma\xi}\tilde{\phi}+\frac{1}{(g^{\prime})^{2}}\partial^{2}_{\xi}\tilde{\phi}+\frac{h^{\prime\prime}}{g^{\prime 2}}\partial_{\sigma}\tilde{\phi}+\left[\frac{1}{2g^{\prime 2}}\mathcal{S}(g)-V_{\ell}\right]\tilde{\phi}=\frac{\mathcal{F}}{\Omega}\,.$ (167)
Starting again from Eq. (145) and introducing the first order (in time) form [see Eq. (147)], we get to the following master wave equation
$\partial_{\sigma}\tilde{\psi}=\frac{1}{1-\left(\frac{h^{\prime}}{g^{\prime}}\right)^{2}}\left\{\frac{2h^{\prime}}{(g^{\prime})^{2}}\partial_{\xi}\tilde{\psi}+\frac{1}{(g^{\prime})^{2}}\partial^{2}_{\xi}\tilde{\phi}+\frac{h^{\prime\prime}}{g^{\prime 2}}\tilde{\psi}+\left[\frac{1}{2g^{\prime 2}}\mathcal{S}(g)-V_{\ell}\right]\tilde{\phi}-\frac{\mathcal{F}}{\Omega}\right\}\,.$ (168)
@@
$\mathcal{L}_{1}\tilde{\phi}$ $=$ $\frac{1}{1-\left(\frac{h^{\prime}}{g^{\prime}}\right)^{2}}\left\{\frac{1}{(g^{\prime})^{2}}\partial^{2}_{\xi}+\left[\frac{1}{2g^{\prime 2}}\mathcal{S}(g)-V_{\ell}\right]\right\}\tilde{\phi}\,,\qquad$ (176)
@@
we observe that the same conformal/Virasoro structure that was present in the Cauchy slices, is reproduced in the hyperboloidal setting.
@@ V Conclusions and Discussion
In this work, we have investigated integrable hidden structures present in the equations governing the dynamics of perturbed BHs. The study reveals an important and novel connection of the dynamics with the properties of integrable systems. From a more global point of view, this work constitutes a step forward towards a longer term objective: the study of structurally stable features in BBH dynamics by adopting an integrability approach based on the effective separation into slow/bulk and fast/boundary DoFs in the description of BBH systems, as proposed in Ref. [6] (see also [7, 8]). In this sense, BHPT offers a perfect arena to start studying some of these questions in a simplified setting in which this separation is clearer. On the other hand, however, it also suggests the possibility of broader applications, not restricted to the dynamics of BBHs. One of the main goals of such program is the study of the interplay between the “bulk” symmetries described here and the asymptotic symmetries at null infinity (the BMS group of symmetries or, possibly, an appropriate extension), an interplay that admits a natural set up in the hyperboloidal approach to BH dynamics  [50]. The present work can then be seen as representing an important and necessary milestone in that direction, by establishing the bulk integrable structures hidden in the dynamics of the slow DoFs.
@@ Acknowledgements.
We warmly thank Badri Krishnan for the shared efforts to define and push the ‘universality from integrability’ program to BBH dynamics. We also thank Carlos Barceló, Jérémy Besson, Piotr Bizoń, Vincent Lam, Corentin Vitel and Mikhail Semenov Tian-Shansky for enlightening discussions. ML and CFS are supported by contracts PID2019-106515GB-I00 and PID2022-137674NB-I00 from MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and 2017-SGR-1469 and 2021-SGR-01529 (AGAUR, Generalitat de Catalunya). ML is also supported by Juan de la Cierva contract FJC2021-047289-I funded by program MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and by NextGenerationEU/PRTR (European Union). This work has also been partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M (Spanish Ministry of Science and Innovation). JLJ is supported by the program "Investissements d’Avenir" through project ISITE-BFC (ANR-15-IDEX-03), the ANR "Quantum Fields interacting with Geometry" (QFG) project (ANR-20-CE40-0018- 02), the EIPHI Graduate School (ANR-17- EURE-0002) and the Spanish FIS2017-86497-C2-1 project (with FEDER contribution).

### v2

@@ II.3 Darboux covariance of BH perturbations
where $\mathcal{C}_{R}$ is an arbitrary integration constant. Notice that logarithmic derivatives of a solution to the time independent Schrödinger equation automatically satisfy the Riccati equation, e.g. $W_{0}$ satisfies it with $\mathcal{C}_{R}=k_{0}^{2}$. The consistency between the expressions for $W$ and $W_{,x}$, as found from Eqs. (39) and (40), gives the following equation for the difference $\delta V=V-v\,$
@@ III Integrable structures in Cauchy slices
We have described the complete landscape of master equations describing BH perturbations, i.e. pairs of master functions and potentials. We have also seen the role of the Darboux transformations as a way of connecting all of them and establishing the isospectrality of the underlying physical description corresponding to this infinite set. In this sense, the Darboux transformation plays the role of a covariance or symmetry of this set of possible physical descriptions of (vacuum) BH perturbations, establishing their physical equivalence. These results provide the correct way to separate the slow and fast DoFs in this context. In this section, we show the hidden integrable structure associated to the nonlinear slow DoFs, i.e. the potentials.  [^8] Note, however, that the integrable structures we identify do not consider the non-radiative $\ell=0,1$ modes, which deserve a separate (future) investigation, possibly involving the need of a refinement/renormalization in the appropriate separation of fast/slow DoFs (we thank the anonymous referee for stressing this point).
@@ III.1.1 Darboux transformation and integrability of the KdV equation
where $\sigma$ is the time of the nonlinear KdV dynamics [^10] Notice that in previous works [15, 16, 17] we used $\tau$ for the KdV evolution instead of $\sigma$. Here we use $\tau$ for the time along the hyperboloidal slicing., which should not be confused with the Schwarzschild time $t$. From the point of view of BH physics, $\sigma$ does not play in principle any role. One interesting analogy with the original KdV equation, which was introduced to study the propagation of waves in shallow water, is that in this context $V$ is essentially the approximation to the gravitational Newtonian potential around the Earth surface, while in our case it is going to be exactly any of the BH potentials (see Sec. II.2).
@@ III.1.2 Isospectrality under KdV flows
The KdV equation, as well as other integrable systems, admits a particular bi-Hamiltonian structure, meaning that one can find two different (but compatible) Poisson structures and Hamiltonians such that the KdV equation can be described in terms of two different by analogous Hamiltonian flows [105]. Starting from the existence of infinite conservation laws [30], Gardner was the first to realize that the KdV equation can be cast as a Hamiltonian system [106]. A different approach was taken by Zakharov and Faddeev [31], who combined inverse scattering techniques together with the Hamiltonian formalism to find the action-angle variables of the system in terms of the Bogoliubov coefficients, which completely defines the scattering matrix in the linear Schrödinger system.
@@ III.2 The structural triangle KdV-Virasoro-Schwarzian derivative
We are rewriting them together to highlight the fact that the fundamental structure responsible for both transformations (the KdV deformations and the conformal transformations) are the same and are given by the Magri bracket. Indeed, this is no coincidence as we can consider the KdV deformations as a restricted family of infinitesimal conformal transformations, as in Eq. (87), by taking
@@ III.3 Conformal transformations in the frequency domain
We are going to restrict ourselves to transformations that cancel the linear term in $\tilde{\psi}_{,\tilde{x}}$, so that the final equation resembles the original Schrödinger equation [Eq. (31)]. By looking into Eq. (102), we see that cancelling the term in $\tilde{\psi}_{,\tilde{x}}$ means to impose the following relation between $\omega$ and $\tilde{x}(x)$:
@@
Therefore, the potential transforms according to the standard conformal transformation of (the holomorphic components of) the energy-momentum tensor of a $2d$ CFT, i.e.
@@
Notice that the transformation of the wave function $\psi$ reminds the conformal transformation, under the mapping $x\mapsto\tilde{x}=\tilde{x}(x)$, of a primary field [114]. Indeed, the KdV/Virasoro structure described in Sec. III.2 relies on the extension to the complex plane (complexification of the independent variable $x$) and on treating the new complex variables $z$ and $\bar{z}$ as independent. The conformal transformation of a primary field in this space is then written as
@@ IV Integrable structures in hyperboloidal slices
We have investigated the emergence of an underlying integrable structure associated to the slow DoFs, i.e. the solitonic potential response of the BH to gravitational perturbations, appearing when studying the linear dynamics of the fast DoFs, the perturbations, in Cauchy slices. As we already mentioned in the Introduction, the separation into slow/fast DoFs shows a natural association to bulk/boundary properties of the system once a hyperboloidal foliation of the Lorentzian manifold $M_{2}$ is considered [54]. In addition, the hyperboloidal slicing is particularly meaningful when studying properties of the system at null infinities. In our framework, it represents a powerful tool to connect bulk and boundary properties of the system. Indeed, we introduce in this section the necessary following step in the construction: the preservation of the integrable structure related to the slow DoFs in the hyperboloidal slicing. This sets the stage for future investigations analyzing the interplay with the boundary/asymptotic features of the system.
In this section we first review the hyperboloidal slicing for master equations and then show that it is invariant under scale transformations of the master function. This feature, of interest by itself, turns out to be crucial in the study of how the bulk symmetries previously described in the Cauchy slices, leak in the hyperboloidal picture, while consistently preserving the bulk/boundary interpretation.
@@ IV.1 Basics of the hyperboloidal slicing
There is an interesting feature of the structure of this operator which provides a different perspective on the separation into slow and fast DoFs. It was shown in Refs. [54, 42] that the non-self-adjointedness of the operator $\mathbb{L}$ is associated to the operator $\mathcal{L}_{2}$, which account for the dissipative character of the system at ${I}^{+}$ and the horizon ($r=r_{s}$), i.e. the asymptotic boundaries. Taking this into account, the operator $\mathcal{L}_{1}$ is associated to “bulk” properties of the system while $\mathcal{L}_{2}$ to “boundary” properties. This supports the conclusion that, within perturbation theory, the bulk properties of the system are expressed by the slow DoFs, in the sense explained above (see Refs. [54, 42] for more detailed arguments).
@@ IV.2 Scale covariance in hyperboloidal slicing
$\mathbb{\tilde{\mathbb{L}}}=\frac{1}{i}\left(\begin{array}[]{cc}0&1\\[5.69054pt] \tilde{\mathcal{L}}_{1}&\tilde{\mathcal{L}}_{2}\end{array}\right)\,,$ (162)
@@
This shows the complete covariance of the hyperboloidal approach under transformations of the master function by a scale factor $\Omega$, as indeed the $\tilde{\mathbb{L}}$ operator preserves the same structure as in the ordinary case (135), with the operator ~$\tilde{\mathcal{L}}_{1}$ having a Sturm-Liouville form. This non-trivial result introduces an apparently arbitrary freedom which may turn out to be useful in several application, including numerical calculations.
@@ IV.3 Conformal hyperboloidal slicing
$-\left[1-\left(\frac{h^{\prime}}{g^{\prime}}\right)^{2}\right]\partial^{2}_{\tau}\tilde{\phi}+\frac{2h^{\prime}}{(g^{\prime})^{2}}\partial^{2}_{\tau\xi}\tilde{\phi}+\frac{1}{(g^{\prime})^{2}}\partial^{2}_{\xi}\tilde{\phi}+\frac{h^{\prime\prime}}{g^{\prime 2}}\partial_{\tau}\tilde{\phi}+\left[\frac{1}{2g^{\prime 2}}\mathcal{S}(g)-V_{\ell}\right]\tilde{\phi}=\frac{\mathcal{F}}{\Omega}\,.$ (167)
Starting again from Eq. (145) and introducing the first order (in time) field $\tilde{\psi}$ [see Eq. (147)], we get to the following master wave equation
$\partial_{\tau}\tilde{\psi}=\frac{1}{1-\left(\frac{h^{\prime}}{g^{\prime}}\right)^{2}}\left\{\frac{2h^{\prime}}{(g^{\prime})^{2}}\partial_{\xi}\tilde{\psi}+\frac{1}{(g^{\prime})^{2}}\partial^{2}_{\xi}\tilde{\phi}+\frac{h^{\prime\prime}}{g^{\prime 2}}\tilde{\psi}+\left[\frac{1}{2g^{\prime 2}}\mathcal{S}(g)-V_{\ell}\right]\tilde{\phi}-\frac{\mathcal{F}}{\Omega}\right\}\,.$ (168)
@@
$\mathcal{L}_{1}\tilde{\phi}=\frac{1}{1-\left(\frac{h^{\prime}}{g^{\prime}}\right)^{2}}\left\{\frac{1}{(g^{\prime})^{2}}\partial^{2}_{\xi}+\left[\frac{1}{2g^{\prime 2}}\mathcal{S}(g)-V_{\ell}\right]\right\}\tilde{\phi}\,,$ (176)
@@
we observe that the same conformal/Virasoro structure that was present in the Cauchy slices, is reproduced in the hyperboloidal setting, namely, the potential transforms as the holomorphic components of the stress-energy momentum tensor.
@@ V Conclusions and Discussion
In this work, we have investigated integrable hidden structures present in the equations governing the dynamics of perturbed BHs. The study reveals an important and novel connection of the dynamics with the properties of integrable systems. From a more global point of view, this work constitutes a step forward towards a longer term objective: the study of structurally stable features in BBH dynamics by adopting an integrability approach based on the effective separation into slow/bulk and fast/boundary DoFs in the description of BBH systems, as proposed in Ref. [6] (see also [7, 8, 9]). In this sense, BHPT offers a perfect arena to start studying some of these questions in a simplified setting in which this separation is clearer. On the other hand, however, it also suggests the possibility of broader applications, not restricted to the dynamics of BBHs. One of the main goals of such program is the study of the interplay between the “bulk” symmetries described here and the asymptotic symmetries at null infinity (the BMS group of symmetries or, possibly, an appropriate extension), an interplay that admits a natural set up in the hyperboloidal approach to BH dynamics  [54]. The present work can then be seen as representing an important and necessary milestone in that direction, by establishing the bulk integrable structures hidden in the dynamics of the slow DoFs.
@@ Acknowledgements.
We warmly thank Badri Krishnan for the shared efforts to define and push the ‘universality from integrability’ program to BBH dynamics. We also thank Carlos Barceló, Jérémy Besson, Piotr Bizoń, Edgar Gasperín, Vincent Lam, Juan A. Valiente-Kroon, Corentin Vitel and Mikhail Semenov Tian-Shansky for enlightening discussions. ML and CFS are supported by contracts PID2019-106515GB-I00 and PID2022-137674NB-I00 from MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and 2017-SGR-1469 and 2021-SGR-01529 (AGAUR, Generalitat de Catalunya). ML is also supported by Juan de la Cierva contract FJC2021-047289-I funded by program MCIN/AEI/10.13039/501100011033 (Spanish Ministry of Science and Innovation) and by NextGenerationEU/PRTR (European Union). This work has also been partially supported by the program Unidad de Excelencia María de Maeztu CEX2020-001058-M (Spanish Ministry of Science and Innovation). JLJ is supported by the program "Investissements d’Avenir" through project ISITE-BFC (ANR-15-IDEX-03), the ANR "Quantum Fields interacting with Geometry" (QFG) project (ANR-20-CE40-0018- 02), the EIPHI Graduate School (ANR-17- EURE-0002) and the Spanish FIS2017-86497-C2-1 project (with FEDER contribution).